A 5.0 L ballon has a pressure of 1.0 atm at sea level, where the temperature is 28°C. The balloon was let loose and ascended to a height in the temperature where the temperature is 11°C. What will be the volume of the balloon if the atmospheric pressure at the new altitude is 0.65 atm?

1
Expert's answer
2022-02-28T11:01:09-0500




"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2},\\\space\\
V_2=V_1\frac{T_2P_1}{T_1P_2}=V_1\frac{(273+11)(1)}{(273+28)(0.65)}=1.45V_1."